# OpenUni: A Simple Baseline for Unified Multimodal Understanding and Generation

### v2

@@ 3.2 Training Recipe
4 Evaluation
Table 3: Results from the GenEval benchmark for text-to-image generation. Here, BLIP3-o-8B* indicates the model that is trained with 30 million additional proprietary data samples. We highlight the best results in bold.
@@ 4 Evaluation
MetaQuery-B [31] - - - - - - 0.74
MetaQuery-L [31] - - - - - - 0.78
MetaQuery-XL [31] - - - - - - 0.80
BLIP3-o-4B [29] - - - - - - 0.81
BLIP3-o-8B [29] - - - - - - 0.83
BLIP3-o-8B* [29] - - - - - - 0.84
OpenUni-B-512 0.99 0.91 0.74 0.90 0.77 0.73 0.84
OpenUni-L-512 0.99 0.91 0.77 0.90 0.75 0.76 0.85
OpenUni-L-1024 0.99 0.92 0.76 0.91 0.82 0.77 0.86
Table 4: Results from the DPG-Bench for text-to-image generation.Here, BLIP3-o-8B* indicates that the model is trained with 30 million additional proprietary data samples. We highlight the best results in bold.

### v3

@@ 3.2 Training Recipe
Table 3: Results from the GenEval benchmark for text-to-image generation. Here, BLIP3-o-8B* indicates the model that is trained with 30 million additional proprietary data samples. $\dagger$ marks results that are obtained by models trained on distillation data (i.e., BLIP3-o-60K [29]). $\ddagger$ denotes rewritten prompts. §$\S$ stands for RL-based methods.
@@
SimpleAR-0.5B-SFT [66] - - - - - - 0.53
SimpleAR-0.5B-RL [66] - - - - - - 0.59§$\S$
SimpleAR-1.5B-SFT [66] - - - - - - 0.61
SimpleAR-1.5B-RL [66] - - - - - - 0.63§$\S$
MetaQuery-B [31] - - - - - - 0.74$\ddagger$
MetaQuery-L [31] - - - - - - 0.78$\ddagger$
MetaQuery-XL [31] - - - - - - 0.80$\ddagger$
BLIP3-o-4B [29] - - - - - - 0.81$\dagger$
BLIP3-o-8B [29] - - - - - - 0.83$\dagger$
BLIP3-o-8B* [29] - - - - - - 0.84$\dagger$
BAGEL [25] 0.99 0.94 0.81 0.88 0.64 0.63 0.82
BAGEL-Rewrite [25] 0.98 0.95 0.84 0.95 0.78 0.77 0.88 $\ddagger$
OpenUni-B-512 0.99 0.91 0.74 0.90 0.77 0.73 0.84$\dagger$
OpenUni-B-512 0.99 0.71 0.55 0.82 0.25 0.42 0.62
OpenUni-L-512 0.99 0.91 0.77 0.90 0.75 0.76 0.85$\dagger$
OpenUni-L-1024 0.99 0.92 0.76 0.91 0.82 0.77 0.86$\dagger$
4 Evaluation
Table 4: Results from the DPG-Bench for text-to-image generation. Here, BLIP3-o-8B* indicates that the model is trained with 30 million additional proprietary data samples. $\P$ denotes the model trained without BLIP3-o-60K [29].
@@ 4.1 Image Generation
Discussion on GenEval. Unlike general text-to-image benchmarks (e.g., DPG-Bench [73]) that use free-form text prompts, GenEval [72] comes with a fixed prompt style: “a photo of A and B”. The prior knowledge of the prompt format has made it much easier to distil GPT4-o [18] or utilize RL methods [63] on this benchmark. In our experiment on OpenUni-B-512, we find utilizing the distillation data (i.e., BLIP3-o-60K [29]) drastically increases the result on GenEval from 0.62 to 0.84 as shown in Table 3. In contrast, we only observe a marginal performance gain on DPG-Bench (79.04 -> 80.29) as shown in Table 4. This has raised concerns over data leakage in the community. Therefore, in Table 3, we labelled all results achieved by distillation data or RL methods, and report our OpenUni’s performance when the distillation data is not available. Additionally, we also mark the results obtained using rewritten prompts.